An object in free fall from

50 m high, so the height is

when t is h(t) = 50 – 4.9t2

.

What is the speed at

t = 2?

2. Determine the equation of the tangent line

on the curve y = x3 at the point (2.8)


1
Expert's answer
2021-11-02T16:12:26-0400

1.

speed:

"v(t)=h'(t)=-9.8t"

"v(2)=9.8\cdot 2=19.6" m/s


2.

"y-y_0=f'(x_0)(x-x_0)"

"f'(x)=3x^2"

"f'(2)=3\cdot 2^2=12"


"y-8=12(x-2)"

"y=12x-16"